OpenStudy (anonymous):

Which trinomial is a perfect square trinomial? (Points : 1) y2 + 6y + 36 y2 + 30y + 60 y2 + 50y + 100 y2 + 24y + 144

OpenStudy (acxbox22):

which one do you think?

rishavraj (rishavraj):

(a + b)^2 = a^2 + b^2 + 2ab

OpenStudy (anonymous):

@Ilovecake @SolomonZelman @Gokuporter @GretaKnows @NeonStrawsForever @pikapow @BlossomCake

OpenStudy (anonymous):

@DDCamp @kool_22 @urbanmorgans @radar @MikeyMaximum @CrazyCountryGirl @Hero

OpenStudy (ilovecake):

No Mass Tag. Thanks

OpenStudy (anonymous):

I don't understand what you're saying. Sorry, could you explain?

OpenStudy (anonymous):

@Ilovecake

OpenStudy (ilovecake):

Tagging a bunch of people in one post. It lags OpenStudy.

OpenStudy (anonymous):

Oh!!! I'm Sorry i didn't know!!

OpenStudy (anonymous):

Could you help though?

OpenStudy (ilovecake):

Yes. :)

OpenStudy (ilovecake):

Do you know what a perfect square trinomial looks like? (a+b)^2 = a^2 + 2(a)(b) + b^2